# Optimal Spatiotemporal Resolutions

### v2

@@ Conclusions
On a practical level, this fundamental problem typically leads to very basic questions, such as: What is the effect of changing the cutoff in analyses conducted using distance-based local descriptors such as SOAP, LENS, coordination numbers, etc.? How to choose the optimal cutoff distance for the study of correlations, transitions, or fluctuations? What is the best inter-frame stride to choose in a time-series analysis to ensure that key physical events are correctly captured? The approach that we report here offers an efficient way to answer such fundamental general questions.
@@
Dealing with the problem in a completely abstract way, our method does not build on any prior assumption and essentially learns the optimal space- and time-resolutions to study a determined system simply based on its data. In this way, the choice of, e.g., the best cutoff distance to be used to translate the system’s trajectories into information-rich time-series data and the best time-interval ($\Delta t$) to extract such information from them, stop being parameters (to be tuned) in the analysis but are implicitly encoded in the physics and in the data of the system itself. The concept of “optimal spatiotemporal resolution” provides a fully unsupervised data-driven approach to identify the optimal spatiotemporal resolutions to characterize any type of complex dynamical system for which a trajectory of the constitutive units is resolvable. The results of the analysis remain thus dependent only on the type of descriptor(s) used (and the type of information that this/these may provide) [martino2024data] and on the resolution in the original acquisition of the raw data (i.e., the frame interval $\Delta\tau$), which sets a maximum in the time-resolution that it is possible to use in the study of the system (i.e., the limit is in the data themselves).
@@
Comparing the results obtained with two very different descriptors for the same system in Fig. 2 vs. Fig. 3 (LENS vs. SOAP), it is interesting to note that any specific descriptor translates the raw data in a different time-series, which has its own optimal spatiotemporal resolution to be best studied. In particular, it is interesting to note that in both cases the optimal spatial resolution to study this aqueous system is found for $r_{c}$ values encompassing up to the 3rd-4th neighbors’ shells (descriptor dependent). At the same time, the optimal time-resolution is different: being a quite noisy local descriptor, SOAP suffers of information loss due to oversampling much more than LENS, which has a higher signal-to-noise ratio [martino2024data, caruso2024classification, lionello2024relevant, crippa_detecting_2023, becchi_layer-by-layer_2024]. While all analyses reported in Figs. 1-5 refer to systems for which it has been possible to resolve the individual trajectories of the constitutive units composing them, it is worth noting that this is not always easy or possible. Nonetheless it is also worth noting that, given the generality of such optimal spatiotemporal resolution concept, this may offer a robust basis to identify optimal resolutions to study and extract information also from different types of data, such as, e.g., static data distributions, or time-series data not derived from many-body systems: in such cases, the broader concept of optimal spatiotemporal resolution simply reduces respectively to that of optimal spatial or temporal resolution.
@@
In general, being the optimal resolution for studying a certain system the one allowing to extract as much information as possible on its internal physics, this automatically unveils the nature of the physical events characterizing the system itself. This is a general principle that is particularly relevant for exploring in unbiased way and for understanding systems whose physics is unknown a priori based on their observation, which is one of the cornerstones of physics. Such a data-driven assessment of “optimal spatiotemporal resolutions” concept provides a simple, yet general and efficient approach for studying and understanding a variety of systems simply based on their data, as well as a practical approach to guide data analysis in general.
@@ Methods
The data shown in Figs 1-3 are obtained from a $50$ ns long MD trajectory of an atomistic model system containing 2048 TIP4P/ICE water molecules [crippa_machine_2023]. The system is simulated at the pressure of 1 atm, starting from a condition in which half of the molecules are in the liquid state and half are arranged in an hexagonal ice packing in correspondence of the melting temperature ($T=268$ K for this force field [conde_high_2017]). Further details on this system can be found in Refs. [crippa_machine_2023, caruso_timesoap_2023, crippa_detecting_2023]. The dynamics of the system is sampled for $\tau=50$ ns of MD simulation with a sampling interval of $\Delta\tau=0.1$ ns.
LENS and SOAP time-series
@@
For each one of the 2048 oxygen atoms, the LENS signals [crippa_machine_2023, crippa_detecting_2023] are computed on the sampled configurations. The different cutoff radii found from the $g(r)$ minima are used to compute different LENS time-series, one for each cutoff ($r_{c}=\{3.3,6,7.5,8.5,10,13,15,20,30\}$).
@@
To describe the structural environment surrounding each particle within the simulations, we used SOAP. We compute the SOAP spectrum, representing the local structural environment of each particle at every timestep within a certain cutoff radius. The SOAP vectors are generated using dynsight platform [noauthor_dynsight_nodate], setting the number of spherical harmonics and number of radial basis functions $n_{\text{max}}=n^{\prime}_{\text{max}}=l_{\text{max}}=8$. The results is a 576-component vector represent the environments of each particle at each timestep. Then, we applied the PCA algorithm to each dataset reducing the dimensionality to the first principal component. SOAP data series have been compute on the system considering the same oxygen atoms and the same cutoff radii as done for LENS.
Onion clustering time-series analyses
@@
For a detailed description of the Onion Clustering method, we refer the interested readers/users to the dedicated paper [becchi_layer-by-layer_2024]. All Onion Clustering analyses conducted in this work used the latest version of the algorithm as implemented and openly available in the dynsight platform [noauthor_dynsight_nodate].
@@
In brief, Onion Clustering is an algorithm for single-point clustering of time-series data, which allows to uncover and discriminate statistically relevant fluctuations from noise and to cluster them based on their similarity [becchi_layer-by-layer_2024, martino2024data, lionello2024relevant, caruso2024classification]. Considering a noisy time-series of total length $\tau$ composed of $n$ time-intervals $\Delta\tau$, Onion Clustering performs a single point clustering of time-series intervals of length set by the time resolution $\Delta t$ used in the analysis. In particular, while the domains classifiable in the time-series depend on the time resolution used, instead of choosing $\Delta t$ a priori Onion Clustering repeats iteratively the analysis from the maximum possible resolution ($\Delta t=2\times\Delta\tau$ in this method) to the minimum one ($\Delta t=\tau$) plotting the obtained results. Looking at the Kernel Density Estimation of the time-series data, the Onion method detects the highest peak corresponding to a first dominant environment, fits a Gaussian distribution to it (i.e., an environment is characterized by the average value of the descriptor, and its variance), and counts the data intervals that remain in that environment for at least $\Delta t$. These data are then classified as part of that environment (stable at that time resolution) and their signals are removed from the time-series that, cleared from their noise, is then analyzed again in search of new high density peaks. This allows to iteratively reduce the noise of the already-classified-data in the time-series and to uncover, layer-by-layer, all microscopic domains that can be statistically classified in robust way at a certain resolution, including those hidden by the noise of the dominant ones. The method iterates until no new statistically relevant clusters can be uncovered. The fraction of data that cannot be classified at every $\Delta t$ is classified as part of the ENV0 cluster (i.e., transitions occurring faster than the time resolution of the analysis). In all the cases studied in this work, Onion Clustering analyses are conducted on the time-series of all particles in the systems.
@@
The data shown in Fig. 4 are obtained from MD simulation described in detail in [cioni_innate_2023, becchi_layer-by-layer_2024, crippa_detecting_2023], of an atomistic model of FCC(211) copper metal parametrized with a deep neural network potential [cioni_innate_2023]. The surface model contains 2400 Cu atoms that are simulated at $T=600$ K. Periodic boundary conditions are applied in the $xy$ plane, while the system is finite along the $z$ direction, simulating de facto an infinite copper/vacuum surface along the (211)-plane. The system is simulated for $\tau=150$ ns, sampling its configuration every $\Delta\tau=10$ ps, for a total of 15000 frames. The LENS signals [crippa_detecting_2023] are then computed for all atoms. The descriptor time-series have been generated for all cutoff in the list $r_{c}=\{3,5,6.5,8.5,10.5,11.5,12.5,14,16\}$. The signals are then smoothed via moving average using a 10 frames window to reduce noise.[crippa_detecting_2023, becchi_layer-by-layer_2024, caruso2024classification]
@@
The data shown in Fig. 5 are obtained by analyzing the trajectories of 6921 Quincke rollers obtained from an experimental microscopy video (reported in Refs. [liu_activity_2021])with a duration of $0.25s$. From the video, the $x$- and $y$-coordinates of 6921 particles for 310 consecutive frames are extracted, using in-house code [becchi_layer-by-layer_2024] and the tracking algorithm Trackpy [allan_trackpy_2016], resulting in a $\Delta\tau=$0.8 ms.
@@
In Eq. 1, the index $j$ runs over the $n_{c}^{i}$ particles located within a specified cutoff distance from particle $i$, and $\vec{v}_{i}$ and $\vec{v}_{j}$ are the velocities of particles $i$ and $j$, respectively. Parameter $\phi_{i}$ ranges from -1 to 1: a value of 1 indicates perfect alignment, meaning particle $i$ and its neighbor particles within $r_{c}$ are moving in the same direction; a value of -1 indicates that velocities are in opposite directions, a value around 0 suggests that the particles are moving in perpendicular directions or exhibit no clear alignment (random relative motion). Denoising was then applied, assigning to each particle at each frame the the value $\Phi_{i}$, equal to the average values of $\phi_{j}$ within the sphere of neighbors. The time-series have been calculate for all cutoff in the list $r_{c}=\{18.9,31.5,46.2,58.8,91,112\}$ $\mu$m.
@@
Complete data and materials pertaining to the trajectories employed, and the analysis conducted herein, are available at https://github.com/GMPavanLab/Optimal-Spatiotemporal-Resolutions (this link will be replaced with a definitive Zenodo link upon acceptance of the final version of this paper). The ”optimal resolution” code used to perform the analyses in this work is available at https://github.com/GMPavanLab/Optimal-Spatiotemporal-Resolutions (and will be also implemented in the dynsight platform https://github.com/GMPavanLab/dynsight shortly). Other information needed is available from the corresponding author upon reasonable request.
Acknowledgments
[…]
@@ Acknowledgments
Figure S1: Supplementary cutoff of Fig 2 from main text. A-B: Onion plots obtained from the analyses of the LENS time-series of water/ice system with $r_{c}=6$(A), $r_{c}=20$(B). The number of resolved micro-clusters is shown in blue, while the fraction of unclassifiable data (in the ENV0 cluster) is shown in orange.
Figure S2: Supplementary cutoff of Fig 3 from main text. A-C: Onion plots obtained from the analyses of the SOAP time-series of water/ice system with $r_{c}=6$(A), $r_{c}=8.5$(B), $r_{c}=20$(C). The number of resolved micro-clusters is shown in blue, while the fraction of unclassifiable data (in the ENV0 cluster) is shown in orange.
Figure S3: Supplementary cutoff of Fig 4 from main text. A-B: Onion plots obtained from the analyses of the LENS time-series of copper system with $r_{c}=11.5$(A), $r_{c}=12.5$(B). The number of resolved micro-clusters is shown in blue, while the fraction of unclassifiable data (in the ENV0 cluster) is shown in orange.
Figure S4: Supplementary cutoff of Fig 5 from main text. A-B: Onion plots obtained from the analyses of the $\Phi$ time-series of Quincke rollers system with $r_{c}=31.5$ $\mu$m (A), $r_{c}=46.2$ $\mu$m (B). The number of resolved micro-clusters is shown in blue, while the fraction of unclassifiable data (in the ENV0 cluster) is shown in orange.

### v4

@@ Conclusions
On a practical level, this fundamental problem typically leads to very basic questions, such as: What is the effect of changing the cutoff in analyses conducted using distance-based local descriptors such as SOAP, LENS, coordination numbers, etc.? How to choose the optimal cutoff distance for the study of correlations, transitions, or fluctuations? What is the best inter-frame stride to choose in a time series analysis to ensure that key physical events are correctly captured? The approach that we report here offers an efficient way to answer such fundamental general questions.
@@
Dealing with the problem in a completely abstract way, our method does not build on any prior assumption and essentially learns the optimal space- and time-resolutions to study a determined system simply based on its data. In this way, the choice of, e.g., the best cutoff distance to be used to translate the system’s trajectories into information-rich time series data and the best time-interval ($\Delta t$) to extract such information from them, stop being parameters (to be tuned) in the analysis but are implicitly encoded in the physics and in the data of the system itself. The concept of “optimal spatiotemporal resolution” provides a fully unsupervised data-driven approach to identify the optimal spatiotemporal resolutions to characterize any type of complex dynamical system for which a trajectory of the constitutive units is resolvable. The results of the analysis remain thus dependent only on the type of descriptor(s) used (and the type of information that this/these may provide) Martino et al. (2024) and on the resolution in the original acquisition of the raw data (i.e., the frame interval $\Delta\tau$), which sets a maximum in the time-resolution that it is possible to use in the study of the system (i.e., the limit is in the data themselves).
@@
Comparing the results obtained with two very different descriptors for the same system in Fig. 2 vs. Fig. 3 (LENS vs. SOAP), it is interesting to note that any specific descriptor translates the raw data in a different time series, which has its own optimal spatiotemporal resolution to be best studied. In particular, it is interesting to note that in both cases the optimal spatial resolution to study this aqueous system is found for $r_{c}$ values encompassing up to the 3rd-4th neighbors’ shells (descriptor dependent). At the same time, the optimal time-resolution is different: being a quite noisy local descriptor, SOAP suffers of information loss due to oversampling much more than LENS, which has a higher signal-to-noise ratio Martino et al. (2024); Caruso et al. (2024); Lionello et al. (2024); Crippa et al. (2023a); Becchi, Fantolino, and Pavan (2024). While all analyses reported in Figs. 1-5 refer to systems for which it has been possible to resolve the individual trajectories of the constitutive units composing them, it is worth noting that this is not always easy or possible. Nonetheless it is also worth noting that, given the generality of such optimal spatiotemporal resolution concept, this may offer a robust basis to identify optimal resolutions to study and extract information also from different types of data, such as, e.g., static data distributions, or time series data not derived from many-body systems: in such cases, the broader concept of optimal spatiotemporal resolution simply reduces respectively to that of optimal spatial or temporal resolution.
@@
In general, being the optimal resolution for studying a certain system the one allowing to extract as much information as possible on its internal physics, this automatically unveils the nature of the physical events characterizing the system itself. This is a general principle that is particularly relevant for exploring in unbiased way and for understanding systems whose physics is unknown a priori based on their observation, which is one of the cornerstones of physics. Such a data-driven assessment of “optimal spatiotemporal resolutions" concept provides a simple, yet general and efficient approach for studying and understanding a variety of systems simply based on their data, as well as a practical approach to guide data analysis in general.
@@ Methods
The data shown in Figs 1-3 were obtained from a $50$ ns long MD trajectory of an atomistic model system containing 2048 TIP4P/ICE water molecules Crippa et al. (2023b). The system was simulated at the pressure of 1 atm, starting from a condition in which half of the molecules are in the liquid state and half were arranged in an hexagonal ice packing in correspondence of the melting temperature ($T=268$ K for this force field Conde, Rovere, and Gallo (2017)). Further details on this system can be found in Refs. Crippa et al. (2023b); Caruso et al. (2023); Crippa et al. (2023a). The dynamics of the system were sampled for $\tau=50$ ns of MD simulation with a sampling interval of $\Delta\tau=0.1$ ns.
LENS and SOAP time series
@@
For each one of the 2048 oxygen atoms, the LENS signals Crippa et al. (2023b, a) were computed on the sampled configurations. The different cutoff radii found from the $g(r)$ minima are used to compute different LENS time series, one for each cutoff ($r_{c}=\{3.3,6,7.5,8.5,10,13,15,20,30\}$).
@@
To describe the structural environment surrounding each particle within the simulations, we used SOAP. We compute the SOAP spectrum, representing the local structural environment of each particle at every timestep within a certain cutoff radius. The SOAP vectors are generated using dynsight platform GMPavanlab (2024), setting the number of spherical harmonics and number of radial basis functions $n_{\text{max}}=n^{\prime}_{\text{max}}=l_{\text{max}}=8$. The results is a 576-component vector represent the environments of each particle at each timestep. Then, we applied the PCA algorithm to each dataset reducing the dimensionality to the first principal component. SOAP data series have been compute on the system considering the same oxygen atoms and the same cutoff radii as done for LENS.
Onion clustering time series analyses
@@
For a detailed description of the Onion Clustering method, we refer the interested readers/users to the dedicated paper Becchi, Fantolino, and Pavan (2024). All Onion Clustering analyses conducted in this work used the latest version of the algorithm as implemented and openly available in the dynsight platform GMPavanlab (2024).
@@
In brief, Onion Clustering is an algorithm for single-point clustering of time series data, which allows to uncover and discriminate statistically relevant fluctuations from noise and to cluster them based on their similarity Becchi, Fantolino, and Pavan (2024); Martino et al. (2024); Lionello et al. (2024); Caruso et al. (2024). Considering a noisy time series of total length $\tau$ composed of $n$ time-intervals $\Delta\tau$, Onion Clustering performs a single point clustering of time series intervals of length set by the time resolution $\Delta t$ used in the analysis. In particular, while the domains classifiable in the time series depend on the time resolution used, instead of choosing $\Delta t$ a priori Onion Clustering repeats iteratively the analysis from the maximum possible resolution ($\Delta t=2\times\Delta\tau$ in this method) to the minimum one ($\Delta t=\tau$) plotting the obtained results. Looking at the Kernel Density Estimation of the time series data, the Onion method detects the highest peak corresponding to a first dominant environment, fits a Gaussian distribution to it (i.e., an environment is characterized by the average value of the descriptor, and its variance), and counts the data intervals that remain in that environment for at least $\Delta t$. These data are then classified as part of that environment (stable at that time resolution) and their signals are removed from the time series that, cleared from their noise, is then analyzed again in search of new high density peaks. This allows to iteratively reduce the noise of the already-classified-data in the time series and to uncover, layer-by-layer, all microscopic domains that can be statistically classified in robust way at a certain resolution, including those hidden by the noise of the dominant ones. The method iterates until no new statistically relevant clusters can be uncovered. The fraction of data that cannot be classified at every $\Delta t$ is classified as part of the ENV0 cluster (i.e., transitions occurring faster than the time resolution of the analysis). In all the cases studied in this work, Onion Clustering analyses are conducted on the time series of all particles in the systems.
@@
The data shown in Fig. 4 were obtained from MD simulation described in detail in Cioni et al. (2023); Becchi, Fantolino, and Pavan (2024); Crippa et al. (2023a), of an atomistic model of FCC(211) copper metal parametrized with a deep neural network potential Cioni et al. (2023). The surface model contains 2400 Cu atoms that were simulated at $T=600$ K. Periodic boundary conditions are applied in the $xy$ plane, while the system was finite along the $z$ direction, simulating de facto an infinite copper/vacuum surface along the (211)-plane. The system is simulated for $\tau=150$ ns, sampling its configuration every $\Delta\tau=10$ ps, for a total of 15000 frames. LENS signals Crippa et al. (2023a) were then computed for all atoms. The descriptor time series have been generated for all cutoff in the list $r_{c}=\{3,5,6.5,8.5,10.5,11.5,12.5,14,16\}$. The signals were then smoothed via moving average using a 10 frames window to reduce noise.Crippa et al. (2023a); Becchi, Fantolino, and Pavan (2024); Caruso et al. (2024)
@@
The data shown in Fig. 5 were obtained by analyzing the trajectories of 6921 Quincke rollers obtained from an experimental microscopy video (reported in Refs. Liu et al. (2021))with a duration of $0.25s$. From the video, the $x$- and $y$-coordinates of 6921 particles for 310 consecutive frames are extracted, using in-house code Becchi, Fantolino, and Pavan (2024) and the tracking algorithm Trackpy Allan et al. (2016), resulting in a $\Delta\tau=$0.8 ms.
@@
In Eq. 1, the index $j$ runs over the $n_{c}^{i}$ particles located within a specified cutoff distance from particle $i$, and $\vec{v}_{i}$ and $\vec{v}_{j}$ are the velocities of particles $i$ and $j$, respectively. Parameter $\phi_{i}$ ranges from -1 to 1: a value of 1 indicates perfect alignment, meaning particle $i$ and its neighbor particles within $r_{c}$ are moving in the same direction; a value of -1 indicates that velocities are in opposite directions, a value around 0 suggests that the particles are moving in perpendicular directions or exhibit no clear alignment (random relative motion). Denoising was then applied, assigning to each particle at each frame the the value $\Phi_{i}$, equal to the average values of $\phi_{j}$ within the sphere of neighbors. The time series were calculate for all cutoff in the list $r_{c}=\{18.9,31.5,46.2,58.8,91,112\}$ $\mu$m.
Acknowledgements.
@@
The data that support the findings of this study are openly available in this Github repository (this link will be replaced with a definitive Zenodo link upon acceptance of the final version of this paper).
Author contributions statement
GMP conceived and supervised this work. DD, SM and MB developed and performed the analyses. All authors discussed the results. DD and GMP wrote the paper. All authors contributed to the revision of the paper.
Conflict of Interest Statement
The authors have no conflicts to disclose.
@@
Figure S1: Supplementary cutoff of Fig 2 from main text. A-B: Onion plots obtained from the analyses of the LENS time series of water/ice system with $r_{c}=6$(A), $r_{c}=20$(B). The number of resolved micro-clusters is shown in blue, while the fraction of unclassifiable data (in the ENV0 cluster) is shown in orange.
Figure S2: Supplementary cutoff of Fig 3 from main text. A-C: Onion plots obtained from the analyses of the SOAP time series of water/ice system with $r_{c}=6$(A), $r_{c}=8.5$(B), $r_{c}=20$(C). The number of resolved micro-clusters is shown in blue, while the fraction of unclassifiable data (in the ENV0 cluster) is shown in orange.
Figure S3: Supplementary cutoff of Fig 4 from main text. A-B: Onion plots obtained from the analyses of the LENS time series of copper system with $r_{c}=11.5$(A), $r_{c}=12.5$(B). The number of resolved micro-clusters is shown in blue, while the fraction of unclassifiable data (in the ENV0 cluster) is shown in orange.
Figure S4: Supplementary cutoff of Fig 5 from main text. A-B: Onion plots obtained from the analyses of the $\Phi$ time series of Quincke rollers system with $r_{c}=31.5$ $\mu$m (A), $r_{c}=46.2$ $\mu$m (B). The number of resolved micro-clusters is shown in blue, while the fraction of unclassifiable data (in the ENV0 cluster) is shown in orange.